OpenStudy (anonymous):

i need help with understanding algebraic equations such as. y^2+y+1/ 2y-2y^4 what would the exclusion be? or how would i find it?

OpenStudy (anonymous):

Well, you don't want your equation to end up looking like this: \[\frac{ stuff }{ 0 }\] That would become an undefined answer/infinity, which is not what this equations is supposed to represent

OpenStudy (anonymous):

so first, find values of y that would make (2y-2y^4)

OpenStudy (anonymous):

then, the exclusion would look something like this: \[y = \mathbb{R}, y \neq yadda\]

OpenStudy (anonymous):

ok. but what is R supposed to stand for? im not familiar

OpenStudy (anonymous):

i got 1+y+y^2,y=0,1-y=0 would this be correct?

OpenStudy (anonymous):

oh, that just means "all real numbers". Basically, any number you can think of. I got y=0 and 1. Where did you get the -1?

OpenStudy (anonymous):

well see its like a multiple choice answer an when i worked it i didnt get the right answer but the closest answer i had to choose was that one so im a little confused. Im working with a program call odyssey ware its online. an hard to teach yourself algebra 2 on

OpenStudy (anonymous):

oh, I can't factor properly :-/ yeah, the answer you placed there is right

OpenStudy (anonymous):

lol ok thank you!!